OpenStudy (anonymous):

simplify the equation 2/2+√3

OpenStudy (jusaquikie):

multiply top and bottom by \[2-\sqrt{3}\] which is called the conjagate

OpenStudy (jusaquikie):

that will remove your radical from the denomonator and make life easier

OpenStudy (anonymous):

@evy15 this process is known as rationalization:)

OpenStudy (anonymous):

ok. working it out and will post my answer in a bit

OpenStudy (jusaquikie):

\[\frac{ 2 }{2+\sqrt{3} }\times \frac{ 2-\sqrt{3} }{2-\sqrt{3} }\]

OpenStudy (jusaquikie):

\[\frac{ 2(2-\sqrt{3}) }{( 2+\sqrt{3})(2-\sqrt{3}) } = \frac{ 4-2\sqrt{3} }{ 4-3} = 4-2\sqrt{3}\]

OpenStudy (anonymous):

yeah this is what I got, thank u so much

OpenStudy (jusaquikie):

np good luck with the rest of your homework.

OpenStudy (anonymous):

thanks